A manufacturer knows that their items have a normally distributed lifespan, with a mean of 3.9 years, and standard deviation of 1 years.

If 10 items are picked at random, 4% of the time their mean life will be less than how many years?

Give your answer to one decimal place.
Answer #1

X :lifespan of items produced by the manufacturer.

X ~ N (3.9 , 1)

if a sample of size(n) 10 is selected, then the sample mean will follow normal distribution with,

mean = 3.9

sd=\frac{\sigma}{\sqrt{n}}=\frac{1}{\sqrt{10}}=0.3162

let , 4% of the time their mean life will be less than m years.

according to the problem,

P(\bar{X}<m)=0.04

\Rightarrow P(\frac{\bar{X}-3.9}{0.3162}<\frac{m-3.9}{0.3162})=0.04

m - 3.9 P(Z < 0.3162) = 0.04

т – 3.9, озво) = 0.04

\Rightarrow \frac{m-3.9}{0.3162}=\Phi^{-1}(0.04)=-1.75069[ in any blank cell of excel type =NORMSINV(0.04) , press enter]

\Rightarrow m=(-1.75069*0.3162)+3.9=3.34643\approx\mathbf{3.3}

If 10 items are picked at random, 4% of the time their mean life will be less than 3.3 years.

***in case of doubt, comment below. And if u liked the solution, please like.